OpenStudy (anonymous):

Medal** Please help Interchange the order of the integration and evaluate the integral

OpenStudy (anonymous):

\[\int\limits_{0}^{pi}\int\limits_{\pi}^{\pi} \frac{ siny }{ y }dydx\]

OpenStudy (astrophysics):

Are those the right intervals?

OpenStudy (anonymous):

yes

OpenStudy (astrophysics):

Ok, think of it this way